OpenStudy (anonymous):

Solve the formula for the given variable. 6x-36=y for x

OpenStudy (anonymous):

y+36)/6

OpenStudy (anonymous):

i put 6 and it says wrong answer

OpenStudy (anonymous):

Add both sides by 36. 6x-36=y 6x-36+36=y+36 6x=y+36 Now, divide both sides by 6. 6x/6 = (y+ 36)/6\[\boxed{x = \dfrac{y+36}{6}}\] Is this clear?